Solve each proportion and then check.
x+10/x-6= x/x+4
HELP ME PLEASE!! :)

Answer by edjones(8007)

(x+10)/(x-6)= x/(x+4) I am guessing that this is the problem. Help tutors. Use parentheses!
(x+10)(x+4)=x(x-6) multiply each side by (x-6)(x+4)
x^2+14x+40=x^2-6x
20x+40=0
20x=-40
x=-2
.
Ed